OpenStudy (anonymous):

FAN AND MEDAL The point (1, −1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (jdoe0001):

\(\begin{array}{cccllll} (1&,&-1)\\ x&,&y \end{array}\qquad hypotenuse=r=\sqrt{x^2+y^2}\impliedby \textit{pythagorean theorem} \\ \quad \\ \quad \\ sin(\theta)=\cfrac{y}{r} \qquad % cosine cos(\theta)=\cfrac{x}{r} \qquad % tangent tan(\theta)=\cfrac{y}{x}\impliedby \textit{find "r", and plug and chug}\)

OpenStudy (anonymous):

Mg I know you from my school :/

OpenStudy (anonymous):

@carsonce I'm kinda homeschooled so...

OpenStudy (danjs):

Draw the thing first

OpenStudy (danjs):

'standard' - relative to the +x axis i believe

OpenStudy (danjs):



OpenStudy (danjs):

4th quadrant

OpenStudy (danjs):

You see how the components of that line are at right angles to the x and y axis?

OpenStudy (danjs):